OpenStudy (anonymous):

how do I solve –2(a + 6) = –22?

OpenStudy (anonymous):

a = 5

OpenStudy (anonymous):

Using the Distributive Property.

OpenStudy (anonymous):

so first you should distribute your -2 to all the terms in the parentheses then get -2a -12 =-22 then add 12 to both sides leaving you with -2a=-10 divide by 2 and get a=5

OpenStudy (anonymous):

-2a-12=-22 +12 +12 _________________ -2a=-10 ____ ____ -2 -2 a=5

OpenStudy (anonymous):

a=5